# LongProLIP: A Probabilistic Vision-Language Model with Long Context Text

### v1

@@ Abstract
Recently, Probabilistic Language-Image Pre-Training (ProLIP) has been proposed to tackle the multiplicity issue of vision-language (VL) tasks. Despite their success in probabilistic representation learning at a scale, the ProLIP models cannot handle long context texts longer than 64 context length, which limits their ability to capture rich contextual information from longer text sequences. To address this issue, this paper proposes a fine-tuning strategy for ProLIP to accept longer texts, e.g., 256 text tokens. Experimental results on Urban-1k and the DataComp evaluation suite show that the proposed LongProLIP recipe can improve understanding of long contexts while minimizing the negative effect of fine-tuning. We also observe a trade-off between the long context understanding (measured by Urban-1k) and general zero-shot capability (measured by ImageNet or the average of 38 zero-shot evaluation datasets by DataComp).
@@ Appendix B Datacomp evaluation suite details
The DataComp evaluate suite contains 38 zero-shot tasks: ImageNet (Russakovsky et al., 2015), 6 ImageNet distribution shifts robust benchmarks, including ImageNet-A, ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), ImageNet v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019) and ObjectNet (Barbu et al., 2019), 13 VTAB task (Zhai et al., 2019), including Caltech-101 (Fei-Fei et al., 2004), CIFAR-100 (Krizhevsky et al., 2009), CLEVR Counts, CLEVR Distance (Johnson et al., 2017), Describable Textures (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), KITTI Vehicle Distance (Geiger et al., 2012), Oxford Flowers-102 (Nilsback & Zisserman, 2008), Oxford-IIIT Pet (Parkhi et al., 2012), PatchCamelyon (Veeling et al., 2018), RESISC45 (Cheng et al., 2017), SVHN (Netzer et al., 2011) and SUN397 (Xiao et al., 2010), and 3 retrieval tasks, including Flickr (Young et al., 2014), MS-COCO Caption (Chen et al., 2015) and WinoGAViL (Bitton et al., 2022). There are also 13 additional tasks, such as CIFAR-10, Country211 (Radford et al., 2021), FGVC Aircraft (Maji et al., 2013), Food-101 (Bossard et al., 2014), GTSRB (Stallkamp et al., 2012), MNIST (LeCun et al., 1998), Pascal VOC (Everingham et al., 2010), Rendered SST2 (Radford et al., 2021), STL-10 (Coates et al., 2011), iWildCam (Beery et al., 2021), FMoW (Christie et al., 2018), Dollar Street (Rojas et al., 2022), and GeoDE (Ramaswamy et al., 2024).

### v2

@@ Abstract
Recently, Probabilistic Language-Image Pre-Training (ProLIP) has been proposed to tackle the multiplicity issue of vision-language (VL) tasks. Despite their success in probabilistic representation learning at a scale, the ProLIP models cannot handle long context texts longer than 64 context length, which limits their ability to capture rich contextual information from longer text sequences. To address this issue, this paper proposes a fine-tuning strategy for ProLIP to accept longer texts, e.g., 256 text tokens. Experimental results on Urban-1k and the DataComp evaluation suite show that the proposed LongProLIP recipe can improve understanding of long contexts while minimizing the negative effect of fine-tuning. We also observe a trade-off between the long context understanding (measured by Urban-1k) and general zero-shot capability (measured by evaluation datasets by DataComp). Code is available at https://github.com/naver-ai/prolip.
@@ Appendix B Datacomp evaluation suite details
The DataComp evaluate suite contains 38 zero-shot tasks: ImageNet (Russakovsky et al., 2015), 6 ImageNet distribution shifts robust benchmarks, including ImageNet-A, ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), ImageNet v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019) and ObjectNet (Barbu et al., 2019), 13 VTAB task (Zhai et al., 2019), including Caltech-101 (Fei-Fei et al., 2004), CIFAR-100 (Krizhevsky & Hinton, 2009), CLEVR Counts, CLEVR Distance (Johnson et al., 2017), Describable Textures (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), KITTI Vehicle Distance (Geiger et al., 2012), Oxford Flowers-102 (Nilsback & Zisserman, 2008), Oxford-IIIT Pet (Parkhi et al., 2012), PatchCamelyon (Veeling et al., 2018), RESISC45 (Cheng et al., 2017), SVHN (Netzer et al., 2011) and SUN397 (Xiao et al., 2010), and 3 retrieval tasks, including Flickr (Young et al., 2014), MS-COCO Caption (Chen et al., 2015) and WinoGAViL (Bitton et al., 2022). There are also 13 additional tasks, such as CIFAR-10, Country211 (Radford et al., 2021), FGVC Aircraft (Maji et al., 2013), Food-101 (Bossard et al., 2014), GTSRB (Stallkamp et al., 2012), MNIST (LeCun et al., 1998), Pascal VOC (Everingham et al., 2010), Rendered SST2 (Radford et al., 2021), STL-10 (Coates et al., 2011), iWildCam (Beery et al., 2021), FMoW (Christie et al., 2018), Dollar Street (Rojas et al., 2022), and GeoDE (Ramaswamy et al., 2024).